# Value Gradients with Action Adaptive Search Trees in Continuous (PO)MDPs

### v1

@@ 3.5 Action Gradient Monte Carlo Tree Search
2:  for all $k=1,\dots,K_{opt}$ do
3:     $g_{a}^{q}=\hat{\nabla}_{a}\hat{Q}(s,a)$Eq. (19) or (17)}
4:     Initialize $a^{acc}\leftarrow a$ for $(s,a)$ if required
5:     aacc
@@
6:     if $\lVert{a^{acc}-a}\rVert>T_{d_{a}}$ then
7:        $addSample\;|\mathbin{=}\;$
@@
PDATE$(s,a,a^{acc})$Eqs. (11), (13); Return TRUE if all $\omega^{i}\leq T_{\omega}^{add}$.}
8:  return $addSample$
We introduce Action Gradient MCTS (AGMCTS), a novel online planning algorithm based on MCTS with DPW. It is applicable to POMDPs via ordered particle belief states, like PFT-DPW [?]. It is described in high-level in Algorithm 1, with novel components highlighted in blue. It combines the MIS estimators for the action-value function, with gradient-based action optimization, and updates actions accordingly. AGMCTS applies action update before expansion. This allows us to pass the information from to determine whether to expand the current action node.
Monte Carlo Gradient Estimation.
@@
While computing (14), (15) can be done based on all children of the current action node, we employ the idea of stochastic gradient optimization to MC estimate the gradient. The first estimate is to sample $K_{O}^{\nabla}$ observation branches with weights $\omega^{i}$. It is possible to uniformly sample and weight by $\omega^{i}$, but we found it to be less effective. The MC approximation of (14) becomes
^$\textstyle\hat{\nabla}_{a}\hat{Q}(s,a)=\frac{1}{K_{O}^{\nabla}}\sum_{k=1}^{K_{O}^{\nabla}}\nabla_{a}\log p_{T}(s^{\prime,i_{k}}\mid s,a)$
$\textstyle\cdot(r(s,a,s^{\prime,i_{k}})+\gamma\hat{V}(s^{\prime,i_{k}}))+\nabla_{a}r(s,a,s^{\prime,i_{k}}).$ (17)
@@
For POMDPs, we estimate (16) with $K_{b}^{\nabla}$ state particles via
$\textstyle\nabla_{a}\log p_{T}(\bar{b}^{-}\mid\bar{b},a)\approx\frac{J}{K_{b}^{\nabla}}\sum_{l=1}^{K_{b}^{\nabla}}\nabla_{a}\log p_{T}(s^{-,j_{l}}\mid s^{j_{l}},a),$ (18)
where indices $j_{l}$ are sampled uniformly from $[1,\dots,J]$. In state-reward POMDPs, we estimate the immediate reward separately. We generate $K_{r}^{\nabla}$ posterior states by sampling $s_{new}^{j}\sim p_{T}(\cdot\mid\bar{b},a)$. The MC estimate of (15) is then
^$\textstyle\hat{\nabla}_{a}\hat{Q}(\bar{b},a)=\frac{1}{K_{r}^{\nabla}}\big{(}\sum_{i=j}^{K_{r}^{\nabla}}\nabla_{a}\log p_{T}(s_{new}^{j}\mid s,a)\cdot$
$\textstyle r(s,a,s_{new}^{j})+\nabla_{a}r(s,a,s_{new}^{j})\big{)}+\big{(}\frac{\gamma}{K_{O}^{\nabla}}\sum_{k=1}^{K_{O}^{\nabla}}(\frac{J}{K_{b}^{\nabla}}\sum_{l=1}^{K_{b}^{\nabla}}$
$\textstyle\nabla_{a}\log p_{T}(s^{-,i_{k},j_{l}}\mid s^{i_{k},j_{l}},a))\textstyle\cdot\hat{V}(\bar{b}^{\prime,i_{k}})\big{)}.$ (19)
Gradient Optimization.

### v3

@@ Appendix B AGMCTS Algorithm
2:  if
A
CTION
O
PT
R
ULE(s,a,d) then
3:     for all $k=1,\dots,K_{opt}$ do
4:        $g_{a}^{q}=\hat{\nabla}_{a}\hat{Q}(s,a)$Eq. (149) or (148)}
5:        Initialize $a^{acc}\leftarrow a$ for $(s,a)$ if required
6:        aacc
@@
7:        if $\lVert{a^{acc}-a}\rVert>T_{d_{a}}^{\min}$ then
8:           $a^{\prime}\;\mathbin{=}\;$
C
LIP
N
ORM$(a^{acc},T_{d_{a}}^{\max})$if $\lVert a^{\mathrm{acc}}\rVert>T_{d_{a}}^{\max}$, scale it down; else keep it}
9:           $addSample\;|\mathbin{=}$
@@
PDATE$(s,a,a^{\prime})$Eqs. (11), (12), (16); Return TRUE if all $\omega_{a^{\prime}}(s^{\prime})\leq T_{\omega}^{add}$.}
10:  return $addSample$
The full AGMCTS algorithm description is outlined in Algorithm 2. It is based on MCTS with DPW, with the new components highlighted in blue. We discuss several of the heuristics and optimizations that we found to be crucial for the performance of AGMCTS.
B.1 Monte Carlo Gradient Estimation
@@
While computing (19), (49) can be done based on all children of the current action node, we borrow from stochastic gradient optimization the idea to use an MC estimate the gradient.
@@
The first MC estimate in AGMCTS is done for POMDPs in the computation of the transition probability log-gradient. We estimate (6) with $K_{b}^{\nabla}$ state particles via
^$\hat{\nabla}_{a}\log p_{T}(\bar{b}^{-}\mid\bar{b},a)=\frac{J}{K_{b}^{\nabla}}\sum_{l=1}^{K_{b}^{\nabla}}\nabla_{a}\log p_{T}(s^{-,j_{l}}\mid s^{j_{l}},a),$ (147)
where indices $j_{l}$ are sampled uniformly from $[1,\dots,J]$. Due to (64), this estimator is unbiased.
In the case we perform exact importance weight update (133), we also sample a number of observation branches to compute the gradient. It is possible to uniformly sample and weight by $\omega_{a}(s^{\prime})$, but we found it to be less effective. The MC approximation of (19) becomes
^$\hat{\nabla}_{a}\hat{Q}(s,a)=\frac{1}{K_{O}^{\nabla}}\sum_{k=1}^{K_{O}^{\nabla}}\nabla_{a}\log p_{T}(s^{\prime,i_{k}}\mid s,a)\cdot(r(s,a,s^{\prime,i_{k}})+\gamma\hat{V}(s^{\prime,i_{k}}))+\nabla_{a}r(s,a,s^{\prime,i_{k}}).$ (148)
In some of the domains we used direct state sampling to compute the immediate reward’s contribution to the gradient based on Theorem 1 Extended (Immediate Reward Gradient), which we found to be more effective. In this form, we generate $K_{r}^{\nabla}$ posterior states by sampling $s_{new}^{j}\sim p_{T}(\cdot\mid\bar{b},a)$. The MC estimate of (49) is then
^$\hat{\nabla}_{a}\hat{Q}(\bar{b},a)=\frac{1}{K_{r}^{\nabla}}\big{(}\sum_{i=j}^{K_{r}^{\nabla}}\nabla_{a}\log p_{T}(s_{new}^{j}\mid s,a)\cdot r(s,a,s_{new}^{j})+\nabla_{a}r(s,a,s_{new}^{j})\big{)}$
$+\big{(}\frac{\gamma}{K_{O}^{\nabla}}\sum_{k=1}^{K_{O}^{\nabla}}(\frac{J}{K_{b}^{\nabla}}\sum_{l=1}^{K_{b}^{\nabla}}\nabla_{a}\log p_{T}(s^{-,i_{k},j_{l}}\mid s^{i_{k},j_{l}},a))\cdot\hat{V}(\bar{b}^{\prime,i_{k}})\big{)}.$ (149)
In the case of linearizing the importance weights update as in equation (12), the summation is over all observation children $s^{\prime}\in C(s,a)$. Since we do not sample them, we weight the summation by the importance weights $\omega_{a}(s^{\prime})$. In this case, the gradient estimator becomes:
^$\hat{\nabla}_{a}\hat{Q}(s,a)=\eta(s,a)^{-1}\sum_{s^{\prime}\in C(s,a)}\omega_{a}(s^{\prime})n(s^{\prime})_{+1}$
$\cdot\left(\nabla_{a}\log p_{T}(s^{\prime}\mid s,a)\cdot(r(s,a,s^{\prime})+\gamma\hat{V}(s^{\prime}))+\nabla_{a}r(s,a,s^{\prime})\right).$ (150)
Similar estimators can be derived for the POMDP case, and for the MDP or POMDP cases when using Theorem 1 Extended (Immediate Reward Gradient).
B.2 Action Update Rule
In some domains, we found that adding a threshold on the number of children $\left|{C(s,a)}\right|\geq K_{\text{child}}^{visits}$ for action update to be performed was beneficial. This improves the estimate of the gradient, since it is the MIS estimator based on the sampled children.
In order to control computational complexity, and also ensure in some domains that enough samples are added to a subtree before performing consecutive action updates, we set a hyperparameter $K_{\text{child}}^{\min}$ for action update to be performed only whenever $n(s,a)\text{ mod }K_{\text{child}}^{\min}=0$.
B.3 Gradient Optimization